Airplane flys at 550mph leaves new york at 9:07am going to california which is 3142mil away. what is the estimated time of arrival? Time on a Clock. using the equation D=v*T D=distance V=valosty T=time
Answer by gonzo(654)

airplane flies for 3142 miles at 550 miles per hour.
R*T=D (rate * time = distance)
T = D/R (time = distance divided by rate
T = time in hours
D = distance = 3142
R = rate = 550 miles per hour
3142/550 = 5.712727273 hours
that equates to 5 hours and .712727273 * 60 minutes
that equates to 5 hours and 42.76 minutes.
that equates to 5 hours and 43 minutes rounded to the nearest minute.
time of departure from new york is 9:07 am new york time (presumed).
plus 5 hours is 2:07 pm new york time.
plus 43 minutes is 2:50 pm new york time.
ETA (estimated time of arrival) is 2:50 pm new york time.
since california time is 3 hours behind new york time,
ETA is 11:50 am california time.